Spectroscopic Model

Best Fit - Synthetic and Experimental Spectra

• Improving atomic structure calculations for high Z neutrals: energy levels and wavefunctions – cross sections and rates

• Reducing the number of calculated cross sections Ag II, Ag III, Ag IV. - multi-representational model.

• Multi-element collisional radiative model - common free e− pool.

• Local plasma effects on the radiator - detailed line shapes

• Non-local effects on the spectra - radiation transport accurate RTE solution.

• Plasma gradients – self-reversal feature.